# Coupling of magnetic and lattice collective excitations in the 2D van der Waals antiferromagnet FePS3

### v1

@@ III.2 Optical Conductivity
The real part of the optical conductivity, $\sigma_{1}$, was extracted from the Drude-Lorenz fit of a slab in a transmission model to the transmission raw data in the range of 40 to 4000  $cm^{-1}$ and is shown in Figure 2. We show the result of our analysis, in which we also employed the Variable Dielectric Function (VDF) method [23] to capture better the details of the raw transmission data at the cryostat’s base temperature of 5 K for both principal polarizations. The same procedure was used at all measured temperatures from 150 K down to 5 K to investigate the temperature dependence of the optical conductivity.
@@
III.3 FePS3 Magnon Excitation
Despite the anisotropy seen in the phononic spectra, the mode at 122  $cm^{-1}$, identified already as the low-energy 15 meV magnon of FePS3 [24, 25, 17], is present in both polarizations with the same spectral weight and at the same frequency. This suggests that the magnon in FePS3 is isotropic compared to the other lattice vibrational modes. Notably, the magnon at 122 $cm^{-1}$ appears in both the Raman shift and the infrared spectra. This observation suggests inversion symmetry breaking in FePS3, a feature not previously reported in the literature [26].
We will now turn to the temperature and magnetic field dependence of the 122  $cm^{-1}$ magnon excitation. As was shown above, the magnon hardens as the temperature is lowered and follows the same temperature dependence for both principal polarizations. Figure 3 shows the magnon’s central frequency as obtained by the Drude-Lorenz fit to our transmission data for both principal axes. To quantitatively describe the temperature dependence of the magnon frequency, we employed a BCS-like order parameter model, as given by the following expression:
@@ III.3 FePS3 Magnon Excitation
where $a$ and $c$ are free parameters of the fit (see caption of Fig. 3 for values) and $T_{N}=118\leavevmode\nobreak\ K$. Eq. 1 (See Eq. 5 in Ref. [27]) is the Bardeen-Cooper-Schrieffer (BCS) type temperature dependence for the superconducting order parameter gap $\Delta$ with critical temperature $T_{c}$, i.e. $\Delta(T)=\Delta(0)\tanh(2.2\sqrt{(T_{c}-T)/T})$ [27, 28, 29]. Here, we show that the magnon central frequency $\omega_{0}$ follows the temperature dependence of a magnetic order parameter with transition temperature $T_{N}$ and provides clear evidence that the 122 $cm^{-1}$ excitation originates at the antiferromagnetic state.
Figure 3: Central frequency of the 122 cm-1 magnon as a function of reduced temperature ($T_{N}/T$) at zero magnetic field. The red line shows a fit to a BCS-like order parameter expression. The inset compares both principal polarizations, indicating isotropic temperature dependence.
Another compelling evidence for the magnetic origin of the 122 $cm^{-1}$ excitation was shown in recent Raman spectroscopy measurements of FePS3 [30] where the excitation, previously considered as a Brillouin-zone folded phonon branch [31, 32], splits into two modes as a function of the magnetic field. Figure 4 shows the field-induced splitting of the 122 $cm^{-1}$ magnon in our transmission spectra, consistent with similar observations in recent optical spectroscopy studies [30, 33]. This splitting becomes clearly resolved at fields up to 7 T. The magnon splitting follows a field dependence with a gyromagnetic ratio of a free electron, and as shown in Figure LABEL:fig:Omega_B for our data in the IR spectra, consistent with previous Raman measurements [30]. However, this field dependence breaks down at higher magnetic fields due to the avoided crossing of two modes, the lower energy magnon branch and the Raman mode at 103  $cm^{-1}$ [33].
Figure 4: Polarized magneto-transmission color maps of FePS3 at 5 K for two principal polarizations, measured as a function of magnetic field (up to 7 T in 1 T steps). Circles mark the central frequencies of the split magnon modes, highlighting the field-induced splitting behavior.
Figure 5: Magnetic-field dependence of the energy splitting $\Delta\nu$ between the central frequencies of the 122 cm-1 magnon mode in FePS3, measured for the two principal polarizations (Pol 1 and Pol 2) at 5 K. The splitting $\Delta\nu$ is extracted from fits to the transmission spectra and plotted as a function of magnetic field $B$. The dashed line represents the expected linear dependence assuming a gyromagnetic ratio ($\gamma\approx 0.94$ cm-1/T), which holds up to $\sim 7$ T before deviations arise due to collective modes’ interactions [33].
@@ III.4 Faraday Angle Rotation
In addition to the infra-red spectroscopy of the temperature and magnetic field dependence of the 122 $cm^{-1}$ magnon excitation, we show in Figure 6 the Faraday angle rotation of FePS3 at 5 K and fields of $B=\pm 7\leavevmode\nobreak\ T$ which was not reported before, to the best of our knowledge. The measurement was obtained using the full protocol for the two different principal axis used in this work and as was explained in the previous section [Sec. II.2].
@@
Figure 6: Faraday rotation angle (top panels) and corresponding magneto-transmission spectra (bottom panels) of FePS3 measured at 5 K under ±$\pm 7$ T magnetic field for two FIR polarizer orientations: (a,c) Pol 1 and (b,d) Pol 2. Pronounced Faraday rotation features appear at the 122 $cm^{-1}$ magnon, showing symmetric splitting with field, as well as at several phonon modes, highlighting strong spin–phonon coupling in the antiferromagnetic state. Transmission below 1% is masked (gray) to avoid unreliable Faraday angle extraction.
@@
The Faraday angle rotation is shown in Fig. 6 along with the transmission as measured by the parallel polarization configuration of the analyzer and polarizer. To avoid artifacts in regions with very low signal, we masked out spectral regions where the transmission coefficient dropped below 1%. This threshold lies above the sensitivity limit of our experimental setup, ensuring the reliability of the Faraday angle measurements. In this way, we can discuss several interesting features, starting from the collective excitations which are originated at the antiferromagnetic spin configuration, i.e. magnons, and following with the lattice collective modes which are coupled to the magnetic order.
[…]
@@
Figure 7: Temperature evolution of the Faraday rotation angle in the 280–350  $cm^{-1}$ range of FePS3, measured at 7 T with FIR polarizer orientation Pol 2. A prominent feature near 320  $cm^{-1}$ emerges below the Néel temperature (TN = 118 K), increases in strength with decreasing temperature, and is attributed to a higher-energy magnon mode. This temperature-dependent behavior further supports the magnetic origin of the excitation.
@@
It is important to note here that the 122 $cm^{-1}$ magnon frequency in our data corresponds to the 15 meV magnon energy detected at the commensurate $q=(0,1,0)$ momentum vector via inelastic neutron scattering measurements [17] with a momentum dispersion reaching about 17 meV at $q=(0,0.5,0)$. On the other hand, the magnon at about 40 meV is seen at the incommensurate momentum of, e.g., $q=(\pm 1,k,\mp 1/3)$ with a momentum dispersion reaching about 17 meV at e.g., $q=(\pm 2,k,\mp 2/3)$ (where $k$ is $0.5,1,1.5$). In general, optical spectroscopy is sensitive only to $q\approx 0$ collective excitations. Nevertheless, the frequencies at which we measure the Faraday angle rotation signal agree with the magnon energies of FePS3 as measured by inelastic neutron scattering. Moreover, it is possible that the difference in intensity between the signal at 122 $cm^{-1}$ (15 meV) and 320 $cm^{-1}$ (40 meV) is associated with the fact that the excitations are commensurate and incommensurate, respectively.
@@ IV Discussion
IV.1 Crystal Symmetries in FePS3
@@
Several vibrational modes of FePS3 exhibit a clear Faraday angle rotation, providing strong evidence for spin–phonon coupling. This observation is consistent with earlier reports [24]. Interestingly, the Raman-active phonon mode at 108 $cm^{-1}$ appears as a weak absorption feature in the infrared magneto-transmission data, where it develops a measurable Faraday rotation of approximately 25 mrad. The same mode is also showing an avoided crossing behavior with the 122 $cm^{-1}$ magnon at high magnetic fields [24].
@@
The appearance of the Raman modes in the IR data suggests a breaking of crystal symmetry. Application of magnetic field naturally breaks time-reversal symmetry. It hence can lead to the mixing of odd and even collective modes, such as those seen in our magneto-transmission measurements but also at the high field magneto transmission data of Ref.  [34], particularly the avoided crossing of the magnon low energy excitation and the Raman mode in the IR data. However, there is no conventional interpretation why this specific Raman mode will also develop a Faraday angle rotation.
@@
Two additional modes show a different behavior, which also cannot be attributed to magnetic field-induced time-reversal symmetry (TRS) breaking. The first mode at 133 $cm^{-1}$ is an infrared active mode but develops a Faraday angle rotation of about 25 mRad. Therefore, this mode already shows a clear spin-phonon coupling through the Faraday angle rotation at low magnetic fields. The second mode at 166 $cm^{-1}$ is also intriguing, as it is a Raman-active mode that appears in the infrared transmission data even at zero magnetic field, but only below the antiferromagnetic transition temperature of FePS3 [30]. This mode also develops a Faraday rotation angle of approximately 50 mrad at 7 T, as shown in Fig. 6.
@@
Figure 8: Color map of polarized mid-infrared transmission at zero field as a function of temperature.
Figure 9: 2D color map of the Raman Intensity in zero field as a function of temperature.
Therefore, the appearance of the 122 $cm^{-1}$magnon, the 166 $cm^{-1}$aman-active phonon, and the mid-infrared excitation at 900 $cm^{-1}$, all visible in both infrared and Raman spectra at zero field, suggests that inversion symmetry is broken in FePS3. This claim is in line with recent experimental indications [35] and in contrast to second harmonic generation measurements [26], which argue that FePS3 preserves inversion symmetry in the magnetic state.
IV.2 Dichroism of the Optical Circular Conductivity
Since we can quantitatively obtain the absolute transmission, the complex conductivity, and the Faraday angle of FePS3 in our measurements, it is, therefore, possible to extract from our data the circular optical conductivity of the split magnon mode. The two field-separated collective modes can be associated with the right- and left-handed circular polarizations of the two opposite spin sublattices in the antiferromagnetic state  [36]. At zero field, there is no difference between one spin population and the other, and the magnon has a single eigen energy. Hence, the two circular polarizations combine into one linearly polarized mode at 122 $cm^{-1}$. It is important to note here that our measurements show no difference between the magnon mode at zero field and the two principal polarizations used in this work. We attribute it to the fact that FePS3 is considered to have an Ising-like (1D) spin Hamiltonian and that the spins are directed along the $c$-axis. Hence, there is no considerable effect on different polarizations along the crystal’s $ab$ plane. However, with the application of a magnetic field, the modes split into low and high energy modes with a corresponding circular optical conductivity $\sigma_{+}$ and $\sigma_{-}$.

### v2

@@ III.2 Optical Conductivity
The real part of the optical conductivity, $\sigma_{1}$, was extracted from the Drude-Lorentz fit of a plan-parallel slab transmission model to the transmission spectra measured in the range of 40 to 4000  $cm^{-1}$  and is shown in Figure 2. We show the result of our analysis, in which we also employed the Variable Dielectric Function (VDF) method [36] to capture better the details of the raw transmission data at the cryostat’s base temperature of 5 K for both principal polarizations. The same procedure was used at all measured temperatures from 150 K down to 5 K to investigate the temperature dependence of the optical conductivity.
@@
III.3 The 122  $cm^{-1}$ Excitation in FePS3
Despite the anisotropy seen in the phononic spectra, the mode at 122  $cm^{-1}$, identified previously as the 122  $cm^{-1}$ magnon of FePS3 [37, 38, 17], is present in both polarizations with the same spectral weight and at the same frequency. This suggests that the 122  $cm^{-1}$ excitation in FePS3 is isotropic compared to the other lattice vibrational modes. Notably, the excitation at 122  $cm^{-1}$ appears in both the Raman shift and the infrared spectra as shown in Fig. 3. This observation suggests inversion symmetry breaking in FePS3, a feature not previously reported in the literature [39].
Figure 3: Comparison of the Raman microscopy spectrum (red) with the transmission spectra for the two principal optical polarizations (black and blue) measured at 5 K and zero magnetic field. The comparison shows the 122 $cm^{-1}$ excitation in all spectra (The negligible slight discrepancy in the Raman shift is acceptable within our systematic error and can be attributed to the calibration of the elastic line in the Raman spectrometer). The figure shows several other excitations discussed in the text, namely the 133 $cm^{-1}$ and 166 $cm^{-1}$ IR modes and the 108 $cm^{-1}$ Raman mode.
We will now turn to the temperature and magnetic field dependence of the 122  $cm^{-1}$ excitation. As was shown above, this mode hardens as the temperature is lowered and follows the same temperature dependence for both principal polarizations. Figure 4 shows the mode’s central frequency as obtained by the Drude-Lorenz fit to our transmission data for both principal axes. Since the temperature dependence resembles that of an order parameter change, we employed the following expression to fit our data:
@@ III.3 The 122  $cm^{-1}$ Excitation in FePS3
where $a$ and $c$ are free parameters of the fit (see caption of Fig. 4 for values) and $T_{N}=118\leavevmode\nobreak\ K$ (See Eq. 5 in Ref. [40] and relevant examples in Ref. [41, 42]). The mode’s central frequency $\omega_{0}$ follows the temperature dependence of a magnetic order parameter with transition temperature $T_{N}$ and provides clear evidence that the 122  $cm^{-1}$ excitation originates at the antiferromagnetic state.
Figure 4: Central frequency of the 122 cm-1 excitation as a function of reduced temperature ($T_{N}/T$) at zero magnetic field. The red line shows a fit to an order parameter temperature dependence (Eq. 1). The inset compares both principal polarizations, indicating isotropic temperature dependence.
Another compelling evidence for the magnetic origin of the 122  $cm^{-1}$ excitation was shown in recent Raman spectroscopy measurements of FePS3 [43] where the excitation, previously considered as a Brillouin-zone folded phonon branch [44, 45], splits into two modes as a function of the magnetic field. Figure 5 shows the field-induced splitting of the 122  $cm^{-1}$ mode in our transmission spectra, consistent with similar observations in recent optical spectroscopy studies [43, 46]. This splitting becomes clearly resolved at fields up to 7 T. The mode’s splitting follows a field dependence with a gyromagnetic ratio of a free electron, and as shown in Figure 6 for our data in the IR spectra, consistent with previous Raman measurements [43]. However, this field dependence breaks down at higher magnetic fields due to the avoided crossing of two modes, the lower energy mode’s branch and the Raman mode at 103  $cm^{-1}$ [46].
Figure 5: Polarized magneto-transmission color maps of FePS3 at 5 K for two principal polarizations, measured as a function of magnetic field (up to 7 T in 1 T steps). Circles mark the central frequencies of the split 122  $cm^{-1}$ mode, highlighting the field-induced splitting behavior.
Figure 6: Magnetic-field dependence of the energy splitting $\Delta\nu$ between the central frequencies of the 122 cm-1 mode in FePS3, measured for the two principal polarizations (Pol 1 and Pol 2) at 5 K. The splitting $\Delta\nu$ is extracted from fits to the transmission spectra and plotted as a function of magnetic field $B$. The dashed line represents the expected linear dependence assuming a gyromagnetic ratio ($\gamma\approx 0.94$ cm-1/T), which holds up to $\sim 7$ T before deviations arise due to collective modes’ interactions [46].
@@ III.4 Faraday Angle Rotation
In addition to the infra-red spectroscopy of the temperature and magnetic field dependence of the 122 $cm^{-1}$ excitation, we show in Figure 7 the Faraday angle rotation of FePS3 at 5 K and fields of $B=\pm 7\leavevmode\nobreak\ T$ which was not reported before, to the best of our knowledge. The measurement was obtained using the full protocol for the two different principal axis used in this work and as was explained in the previous section [Sec. II.2].
@@
Figure 7: Faraday rotation angle (top panels) and corresponding magneto-transmission spectra (bottom panels) of FePS3 measured at 5 K under $\pm 7$ T magnetic field for two FIR polarizer orientations: (a,c) Pol 1 and (b,d) Pol 2. Pronounced Faraday rotation features appear at the 122 $cm^{-1}$ excitation, showing symmetric splitting with field, as well as at several phonon modes, highlighting strong spin–phonon coupling in the antiferromagnetic state. Transmission below 1% is masked (gray) to avoid unreliable Faraday angle extraction.
@@
The Faraday angle rotation is shown in Fig. 7 along with the transmission as measured by the parallel polarization configuration of the analyzer and polarizer. To avoid artifacts in regions with very low signal, we masked out spectral regions where the transmission coefficient dropped below 1%. This threshold lies above the sensitivity limit of our experimental setup, ensuring the reliability of the Faraday angle measurements. In this way, we can discuss several interesting features, starting from the collective excitations that originate at the antiferromagnetic state and then by the lattice collective modes that are coupled to the magnetic order.
The 122 $cm^{-1}$ excitation that splits into two modes shows a Faraday angle rotation of about 0.1 to 0.15 rad at 5 K and 7 T. This Faraday angle is much larger than the sensitivity of the measurements setup (about 2 mRad) and well above the measurement noise level, as shown in Figure 7. In addition to the 122 $cm^{-1}$  it is known that FePS3 has another excitation at higher energies of about 40 meV. Close examination of our Faraday angle rotation spectrum close to that energy reveals a very weak feature, which we attribute to this high-energy excitation. Figure 8 shows the Faraday rotation angle for selected temperatures above and below the antiferromagnetic transition of FePS3. Below $T_{N}$ but still close to the phase transition, it is very hard to detect any Faraday angle rotation above our measurement baseline and within the noise level however, as the system goes deep into the antiferromagnetic phase, a dip in the data appears at about 320 $cm^{-1}$ which corresponds to an energy of about 40 meV. The appearance of a Faraday rotation signal at $\leavevmode\nobreak\ 320$ $cm^{-1}$, combined with a temperature-dependent shift and intensity increase, provides strong evidence that this feature originates from a magnetic excitation.
@@
Figure 8: Temperature evolution of the Faraday rotation angle in the 280–350  $cm^{-1}$ range of FePS3, measured at 7 T with FIR polarizer orientation Pol 2. A prominent feature near 320  $cm^{-1}$ emerges below the Néel temperature (TN = 118 K), increases in strength with decreasing temperature, and is attributed to a higher-energy mode at 40 meV. This temperature-dependent behavior further supports the magnetic origin of the excitation.
@@
It is important to note here that the frequency 122 $cm^{-1}$ in our data corresponds to the energy of 15 meV detected at the commensurate $q=(0,1,0)$ momentum vector via inelastic neutron scattering measurements [17] with a momentum dispersion reaching about 17 meV at $q=(0,0.5,0)$. On the other hand, the excitation at about 40 meV is seen at the incommensurate momentum of, e.g., $q=(\pm 1,k,\mp 1/3)$ with a momentum dispersion reaching about 17 meV at e.g., $q=(\pm 2,k,\mp 2/3)$ (where $k$ is $0.5,1,1.5$). In general, optical spectroscopy is sensitive only to $q\approx 0$ collective excitations. Nevertheless, the frequencies at which we measure the Faraday angle rotation signal agree with the excitation energies of FePS3 as measured by inelastic neutron scattering. Moreover, it is possible that the difference in intensity between the signal at 122 $cm^{-1}$ (15 meV) and 320 $cm^{-1}$ (40 meV) is associated with the fact that the excitations are commensurate and incommensurate, respectively.
@@ IV Discussion
IV.1 Electronic structure of FePS3
Figure 9: Local crystal field environment of Fe2+ in FePS3 and orientation of the $e_{g}$ (top) and $t_{2g}$ (bottom) orbitals. The $e_{g}$ states have one protuberance pointing toward the negative sulfur ions and one away from it. The $t_{2g}$ states point away from the negative ions. If we call the axis connecting the farthest corners of the two cubes the $z$ axis, the yzyz and zxzx orbitals have one protuberance along $z$ and one protuberance in the plane. The xyxy orbital has both protuberances pointing in the plane (i.e. away from the negative sulfur ions).
Fe2+, has a $3d^{6}$ configuration. The 210 possible $3d^{6}$ states are described by strongly correlated many-electron wavefunctions as a result of the Coulomb interaction. Thanks to the symmetries of the problem, these states can be labeled according to spin, orbital momentum, and total (spin plus orbital) angular momentum of the interacting electron system. In the absence of a crystal field and spin-orbit interaction, the ground state is described by the $^{5}D$ multiplet comprising $5\times 5$ degenerate eigenstates. The remaining 185 eigenstates are grouped in multiplets scattered over a range of 10 eV above the ground state [47].
In FePS3 the negatively charged sulfur ions surrounding Fe2+ occupy roughly the 8 corners of an octahedron (see Fig. 9). The ensuing crystal field splits the $^{5}D$ multiplet in $^{5}T_{2g}$ and $^{5}E_{g}$. The $^{5}E_{g}$ states have protuberances closer to the negative sulfur ions than the $^{5}T_{2g}$. Consequently, the energy of the $^{5}T_{2g}$ states is about 1 eV lower than that of the $^{5}E_{g}$ states.
Spin-orbit coupling splits the $^{5}T_{2g}$ states in three multiplets (see Fig. 10): The ground state has $J=1$, an excited $J=3$ multiplet at $\sim 30$ meV and $J=5$ at $\sim 60$ meV. A further refinement comes from the trigonal crystal structure, namely that two of the $^{5}T_{2g}$ states have one protuberance in the plane (i.e. away from the negative sulfur ions) and one perpendicular to it. The remaining $^{5}T_{2g}$ state has both protuberances pointing in the plane (i.e. away from the negative sulfur ions). The trigonal field $\Delta$ is then negative and separates the $^{5}T_{2g}$ multiplet into a $^{5}A_{g}$ ground state and a $^{5}E_{g}$ excited state. Refs. [23, 48, 49, 13, 50, 51, 52] also obtained/assumed $\Delta<0$. On the other hand, a positive sign was obtained in Ref. [53] by fitting high-temperature susceptibility data, see Fig. 10.
Figure 10: Tanabe-Sugano diagram of the $d^{6}$ energy levels in a trigonal crystal field adapted from Ref. [53]. The black arrow on the right is from the original paper. The green arrow on the left is our tentative assignment.
Although the orbital momentum is quenched by the crystal field, a finite orbital moment is induced by the spin-orbit coupling. This is illustrated in Fig. 10, which was adapted from Ref. [53]. The arrow on the right-hand panel is from the original paper, but in view of the expected sign of the crystal field discussed above, this seems unphysical. With this sign of crystal field the ground state of an isolated Fe2+ ion has no magnetic moment, whereas there is a moment for $\Delta<0$. For $\Delta>0$, the possibility remains that a magnetic moment is induced by the Weiss field [53].
The energy levels of our tentative assignment (green arrow in Fig. 10, $\Delta<0)$ are reproduced in Table 2. The highest energy states are consistent with observed absorption features in the range of 1000 cm-1. It is important to strike a note of caution at this point: the model used to generate this table did not take into account the orbital mixing of Fe $3d$ and the ligand states [54], and this cannot be remedied by a rescaling of interaction and/or crystal field parameters because the degree of orbital mixing will in general be different from one multiplet to another according to its shape and orientation. Consequently, the energy spacings in this Table 2 cannot be fully accurate.
Energy (meV) Degeneracy
126.6 2
126.5 1
120.2 2
61.4 2
41.4 2
26.9 2
17.9 1
6.4 1
0 2
Table 2: Energies of the $^{5}T_{2g}-$manifold split by crystal field and spin-orbit parameter, corresponding to the yellow arrow in the left panel of Fig. 10
The ground state is a spin-orbit doublet with a finite magnetic moment. In an AF environment the exchange field of the surrounding ions lifts the degeneracy of the ground state, making the spin point up or down. Applying a magnetic field along $z$ has the effect for half of the spins that their energy increases, and for the other half it decreases. If one now observes the transition from the ground state to the singly degenerate excited state at 17.9 meV, the spectrum will show a splitting of this transition, the peak moving up coming from the down-spin atoms, and the peak moving down coming from the up-spin atoms.
Another potential interesting feature is the transition from the ground state to the singly degenerate 6.4 meV excited state. Formally this can be identified with the anti-ferromagnetic resonance [55]. In Ref.  [55] the phenomenon is described as the precession of the spin in an anisotropy field $H_{a}$ with frequency $\hbar\omega/\gamma=H_{0}\pm[H_{a}(2H_{e}+H_{a})]^{1/2}$ where $H_{0}$ is the external field, $\gamma=ge/2mc$ and $H_{e}$ is the Weiss exchange field. In the local impurity case $H_{a}$ is replaced by the splitting of the three-fold degenerate spin-orbital state in a ground state doublet and an excited singlet state. It is important to emphasize that what we usually call the “magnon” in an optical spectrum is the lowest excitation, in the present case, 6.4 meV. The state at 17.9 meV constitutes a rearrangement of the many-body wave-function compared to the ground state, and is best referred to as a $dd$-exciton.
IV.2 Crystal Symmetries in FePS3
@@
We previously showed in Fig. 3 a comparison between our polarized transmission data and Raman microscopy done on similar FePS3 of the same growth batch. Several vibrational modes of FePS3 exhibit a clear Faraday angle rotation, providing strong evidence for spin–phonon coupling. This observation is consistent with earlier reports [37]. Interestingly, the Raman-active phonon mode at 108 $cm^{-1}$ appears as a weak absorption feature in the infrared magneto-transmission data (See Fig.  3 and Table 1), where it also develops a measurable Faraday rotation of approximately 25 mrad. The same mode is also showing an avoided crossing behavior with the 122 $cm^{-1}$ excitation at high magnetic fields [37].
@@
The appearance of the Raman modes in the IR data suggests a breaking of inversion symmetry and the mixing of odd and even collective modes. The application of a magnetic field naturally breaks time-reversal symmetry and can result in the appearance of Raman excitations in our magneto-transmission measurements. This observation is also shown in the high field magneto transmission data of Ref.  [56], particularly the avoided crossing of the 122 $cm^{-1}$ excitation and the Raman mode in the IR data. However, there is no conventional interpretation of why this specific Raman mode will also develop a Faraday angle rotation.
@@
Two additional modes show a different behavior, which also cannot be attributed to magnetic field-induced time-reversal symmetry (TRS) breaking. The modes at 133 $cm^{-1}$ and 166 $cm^{-1}$ are infrared active modes (see comparison with the Raman data in Fig. 3 and Table 1) but develop a Faraday angle rotation of about 25 mRad and 50 mrad, respectively, at 7 T as seen in Fig. 7. Therefore, these modes already show a clear spin-phonon coupling through the Faraday angle rotation at low magnetic fields. It is possible that the phonon symmetry in this case is coupled to the in-plane anisotropy of the magnetic order, being different along the $a$-axis and along the $b$-axis, resulting in magnetic field-induced chirality of the phonons’ optical response [57].
@@
Figure 11: Top panel: Color map of polarized infrared transmission (from 50  $cm^{-1}$and for Pol 1) at zero field as a function of temperature. Bottom panel: 2D color map of the unpolarized Raman Intensity in zero field as a function of temperature. The temperature range in both measurements was obtained between 10 K to 150 K in intervals of 10 K.
Therefore, the appearance of the 122 $cm^{-1}$ excitation and the mid-infrared excitation at 900 $cm^{-1}$, all visible in both infrared and Raman spectra at zero field, suggests that inversion symmetry is broken in FePS3. This claim is in line with recent experimental indications [58] and in contrast to second harmonic generation measurements [39], which argue that FePS3 preserves inversion symmetry in the magnetic state.
IV.3 Dichroism of the Optical Circular Conductivity
FePS3 is considered to have an Ising-like (1D) spin Hamiltonian where the spins are directed along the $c$-axis. Hence, there should be a negligible, if any, anisotropy effect on the absorption spectra in different polarizations along the crystal’s $ab$ plane, as can be seen in our data (see App. A). This allows us to consider the 122 $cm^{-1}$ as an isotropic excitation in the polarized magneto-transmission data. Since we can quantitatively obtain the absolute transmission, the complex conductivity, and the Faraday angle of FePS3 in our measurements, it is, therefore, possible to extract from our data the circular optical conductivity of the split 122 $cm^{-1}$ mode. The two field-separated collective modes can be associated with the right- and left-handed circular polarizations of the two opposite spin sublattices in the antiferromagnetic state  [59]. At zero field, there is no difference between one spin population and the other, and the 122 $cm^{-1}$ excitation has a single eigen energy. Hence, the two circular polarizations combine into one linearly polarized mode at 122 $cm^{-1}$. The application of a magnetic field splits the modes into low and high energy modes with a corresponding circular optical conductivity $\sigma_{+}$ and $\sigma_{-}$.